OpenStudy (anonymous):

(2u^2-u+7)+(-4u^2+6u+5)

OpenStudy (anonymous):

\[\Delta = 121\]

OpenStudy (jango_in_dtown):

(2u^2-u+7)+(-4u^2+6u+5)=(2-4)u^2+(-1+6)u+(7+5)=-2u^2+5u+12

OpenStudy (anonymous):

i got -2u^2+5u+12..... okay thank you. wanted to be sure

OpenStudy (anonymous):

can you check another one for me?

OpenStudy (anonymous):

o u wanted the result xD sry was confuzed on that part

OpenStudy (anonymous):

yeah

OpenStudy (jango_in_dtown):

ok

OpenStudy (anonymous):

9z^4(4-6z^2)

OpenStudy (anonymous):

i got 36z^4-54z^6 but i have to simplify it as much as possible

OpenStudy (jango_in_dtown):

Its correct

OpenStudy (anonymous):

okay thank you